bill533:

What is the closed linear form for this sequence given a1 = 14 and an + 1 = an - 2? A) an = 16 + 2n B) an = 16 - 2n C) an = 14 + 2n D) an = 14 - 2n E) an = 16 - n2

bill533:

B

bill533:

@Elsa213

Elsa213:

@AngeI o:

AngeI:

What kind of math is this o.O

bill533:

Functions

AngeI:

ew Have fun Dude

dude:

Agh notation I'm going to assume that it is \(a_1 = 14\) and \(a_{n + 1} = a_n - 2\) \(a_1 = 14\) would be point (1,14) \(a_{n + 1} = a_n - 2\) indicates its rate "-2x" y=-2x+b Since we know a point (1,14), just substitute 14=-2(1)+b b=16 Put it into the equation y=-2x+16 or just \(a_n=-2n+16\) or \(a_n = 16 - 2n\) Yes it's right